# Thermodynamic Basis of Sugar-Dependent Polymer Stabilization: Informing Biologic Formulation Design

### v1

@@ 2.1 System Setup and Molecular Dynamics Simulations
We simulated a united-atom HP model with 26 monomeric units, where each monomer represents a methylene ($\mathrm{CH_{2}}$) unit in terms of size and mass. The Lennard-Jones parameters used for the polymer model are the same as those used previously to represent $\mathrm{CH_{2}}$.64, 65, 58 ($\sigma=0.373$ $\mathrm{nm}$ and $\epsilon=0.5856$ $\mathrm{kJ/mol}$). The CP model is similar to the HP model with additional partial charges ($\pm 0.5$ e) assigned to 4 beads (Fig. 1a). The net charge of the CP model is zero, ensuring that any observed differences arise from local electrostatic effects rather than overall charge imbalance. Replica exchange umbrella sampling (REUS) simulations (see Section 2.2) were performed using GROMACS 2021.7 with the PLUMED 2.9.0 patch applied. Kirkwood-Buff derived parameters (KBP) developed by Cloutier et al. 66 were used for the sugars ($\alpha$-glucose, $\beta$-fructose, trehalose, sucrose) considered in this study (Fig. 1b). This force field more accurately describes protein-excipient-water interactions than standard force fields.
We employed the TIP4P/2005 water model67 to describe water. TIP4P/2005 has been shown to accurately predict several water properties such as density, viscosity, diffusion coefficients, and the hydrogen-bonding network 68, 67. Since the original KBP were developed for use with the TIP3P water model,69 we assessed their applicability to the TIP4P/2005 water model. We calculated the Kirkwood-Buff integral for the sugar-water systems and compared the results with those reported by Cloutier et al.66. The values obtained are within the range of experimental error, confirming the compatibility of KBP with TIP4P/2005 model (see Table S1 for details). We performed MD simulations of HP and CP models in sugar solutions of concentrations ranging from 0 to 2.0 M (see Table S2 for further details). Monosaccharide concentrations were studied up to 2.0 M and disaccharide concentrations were limited to 1.0 M to ensure an equivalent monomeric unit concentration.
@@
The systems in all simulations were first energy minimized using the steepest descent algorithm.70 NVT equilibration was carried out for 1 ns at 300 K, followed by a 1 ns NPT equilibration at 300 K and 1 atm. During equilibration, temperature was controlled with the V-rescale thermostat,71 and pressure was controlled with the Berendsen barostat.72 Following equilibration, NPT production runs were completed using the Nosé-Hoover thermostat73 and Parrinello-Rahman barostat.74 Production runs were completed for 100 ns per window for polymer/sugar/water REUS simulations. 100 ns was sufficient to achieve convergence, as demonstrated by the evolution of the PMF with simulation time (SI Fig. S1). The Particle Mesh Ewald (PME) algorithm was used for long-range electrostatic interactions with a cut-off of 1 $\mathrm{nm}$.75 A reciprocal grid of 42 x 42 x 42 cells was used with $\mathrm{4^{th}}$ order B-spline interpolation. A single cut-off of 1 $\mathrm{nm}$ was used for van der Waals interactions. The neighbor search was performed every 10 steps. Lorentz-Berthelot mixing rules76, 77 were used to calculate non-bonded interactions between different atom types, except for polymer-water oxygen interactions. The polymer-water oxygen interactions were adjusted such that the folded and unfolded states of the polymer in pure water were approximately evenly distributed in straightforward MD simulations.64, 65
@@ 2.2 Replica Exchange Umbrella Sampling
REUS was employed to sample the conformational free energy landscape of the polymer in different sugar solutions and obtain the potential of mean force (PMF) as a function of the polymer radius of gyration ($R_{g}$). 12 evenly spaced umbrella potential window centers from 0.35 to 0.90 nm were used, with a spacing of 0.05 nm between consecutive centers. A harmonic bias potential of 5000 kJ/mol/nm2 was applied to all the windows except the window at 0.45 nm, where 1000 kJ/mol/nm2 was applied. This choice of force constants was adopted from our previous work on a similar system, where it ensured adequate sampling and convergence of the free energy profile.64, 65
@@
where $R_{g,cut}$ is determined as the transition point between the folded and unfolded states where $\frac{dW(R_{g})}{dR_{g}}=0$. $W(R_{g})$ is the PMF, calculated as $W(R_{g})=$} $k_{B}T\ln(P(R_{g}))$}. The Weighted Histogram Analysis Method (WHAM) was used to reweight biased probability distributions obtained from REUS simulations.78
@@ 2.3 Thermodynamic Decomposition
We decomposed the PMF into thermodynamic components to further investigate the role of sugars in HP and CP folding. Following the methods established in prior studies,59, 79, 80 the PMF can be separated into two contributions – intrapolymer interactions and solvent-mediated interactions ($W_{solv}$). The former can be estimated from simulations of polymer in vacuum ($W_{vac}$), leading to the expression:
@@
Consistent with the solvation theory approaches, 81, 82, 83, 84, 85, 86, 87, 88, 42, 89 $\Delta G_{solv}$ can be conceptually decomposed into two steps: (1) the reversible work required to create a polymer-sized cavity in the solvent, and (2) the energy contributions arising from "turning on" polymer-solvent interactions upon solvation. This relationship was expressed by Arsiccio et al.42 as:
@@
Here, $\Delta G_{ex}$ represents the free energy associated with creating a cavity in the solvent (also known as the excluded volume contribution), and $\Delta G_{si}$ represents the contribution from polymer-solvent interactions (also known as the soft-interaction contribution). The sum of enthalpic components, $\Delta H_{ex}$ + $\Delta H_{si}$, can be approximated as the polymer-solvent interaction energies under the assumption that pressure-volume (PV) work is negligible.
@@ 2.4 Preferential Interaction Coefficient
where $\mu$ is the chemical potential, $m$ is the concentration, subscripts $p$, and $s$ refer to polymer and sugar, respectively, and T and P refer to the temperature and pressure. This parameter is calculated in simulations using the two-domain formula36, 90, 91 given by:
@@ 3.1 Effect of Sugars on Polymer Folding
The free energy of unfolding provides a measure of the stability of the ensemble of the folded vs. unfolded states of the polymer. For monosaccharides, the free energy of unfolding increases as the sugar concentration increases up to 1 M. This is consistent with previous observations using the linear extrapolation method, which assumes that the free energy of unfolding varies monotonically with cosolvent concentrations and is commonly used to estimate stability changes in the presence of denaturants or stabilizers. 92, 93, 42 However, at 2 M concentration, the free energy of unfolding decreases, suggesting a non-monotonic, concentration-dependent influence of sugars on polymer stability. This shift occurred at lower concentrations for disaccharides (0.5 M) than monosaccharides, consistent with the trends observed in the unfolded state region of the PMF.
@@ 3.2 Role of Preferential Exclusion on Polymer Stabilization
We compute the preferential interaction coefficient ($\Gamma$) for the HP and CP models in different sugar solutions. Our results consistently show positive $\Gamma$ values across all sugar solutions, indicating preferential accumulation of sugars around HP and CP. The magnitude of $\Gamma$ increases with sugar concentration, with disaccharides exhibiting higher values than monosaccharides at the same molar concentrations. On a mass basis, disaccharides still tend to exhibit higher $\Gamma$ values than monosaccharides, with the notable exception of 2 M $\beta$-fructose, which shows values comparable to 1 M sucrose.
The preferential exclusion mechanism 43, 42 suggests that stabilizing sugars are preferentially excluded from (folded) proteins, resulting in negative $\Gamma$ values. However, our results show positive $\Gamma$ values even in cases where the folded states of HP and CP are thermodynamically favored, seemingly contradicting this mechanism. This apparent inconsistency can be reconciled by recognizing that preferential exclusion is fundamentally governed not by the absolute magnitude of $\Gamma$, but by the difference in $\Gamma$ between the folded and unfolded states. Specifically, the mechanism assumes greater exclusion from the unfolded state, and it is this difference in $\Gamma$ that contributes to the stabilization. To capture this effect, we examine the difference in preferential interactions: $\Delta\Gamma=\Gamma_{folded}-\Gamma_{unfolded}$. At lower sugar concentrations, $\Delta\Gamma$ is positive, indicating stronger sugar accumulation around the folded state. This correlates with an increase in unfolding free energy with concentration, suggesting increasing stabilization of the folded state. At higher concentrations, $\Delta\Gamma$ becomes negative, and this transition coincides with a decrease in unfolding free energy with increasing sugar concentration. This trend is consistent with the Wyman linkage relation 94, which links changes in solute binding to changes in macromolecular stability. Wyman linkage relation for folding/unfolding equilibria can be expressed as:94
@@ 3.3 Thermodynamic Decomposition of Unfolding Free Energy
Figure 5: Decomposition of HP unfolding free energy contributions in (a) $\alpha$-glucose, (b) $\beta$-fructose, (c) trehalose, and (d) sucrose solutions. Positive values increase the stability of the folded state and negative values destabilize the folded state of the HP. Increasing additive concentration is denoted by increased shading (light to dark).
Figure 6: Decomposition of CP unfolding free energy contributions in (a) $\alpha$-glucose, (b) $\beta$-fructose, (c) trehalose, and (d) sucrose solutions. Positive values increase the stability of the folded state and negative values destabilize the folded state of the CP. Increasing additive concentration is denoted by increased shading (light to dark).
@@
For HP in monosaccharide solutions (Fig. 5a, b), polymer-sugar interactions ($\Delta\Delta E_{ps}$) stabilize the folded state up to a concentration of 0.5 M. At 1 M, this effect reduces, and by 2 M, these interactions favor the unfolded state. This transition suggests that at higher sugar concentrations, additional binding sites on the polymer that become accessible upon unfolding interact preferentially with sugars, thereby stabilizing the unfolded conformation. This is also largely supported by the increased $\Gamma$ values for the unfolded state at high sugar concentrations. These findings explain the decrease in $\Delta\Delta G_{u}$ observed at high sugar concentrations. Conversely, polymer-water interactions ($\Delta\Delta E_{pw}$) favor unfolding at concentrations up to 0.5 M but promote folding at higher concentrations. $\Delta\Delta G_{rem}$, representing the entropic effects, consistently favors unfolding in all sugar solutions relative to pure water. In disaccharide solutions (Fig. 5c, d), a similar trend is observed, but the stabilizing effect of $\Delta\Delta E_{ps}$ shifts toward favoring the unfolded state as early as 0.5 M. This earlier onset of stabilization compared to monosaccharides can be attributed to the larger size of disaccharides: having one disaccharide molecule in solution is approximately equivalent, in mass or volume, to having two monosaccharide molecules. As a result, trends that appear at 1 M for monosaccharides tend to emerge at roughly 0.5 M in disaccharide solutions. This effective scaling helps explain why $\Delta G_{u}$ decreases at lower concentrations in disaccharides relative to monosaccharides. While this mass-based scaling captures broad concentration-dependent trends, deviations between sugars at equal mass concentrations indicate that other molecular features also contribute to the observed behavior, explained further in Section 3.4 and 3.5.
@@
As discussed in Section 3.3, $\Delta\Delta G_{rem}$ consists of entropic effects from cavity creation (excluded volume contribution) and polymer-solvent interactions (referred to as the interaction entropy contribution). Previous studies suggest that the excluded volume contribution is directly related to solvent surface tension. 43, 49, 42 Specifically, higher surface tension increases the energy required for cavity formation, favoring smaller cavities and thus stabilizing the folded state. Since all sugars considered here increase the air-water surface tension 43, 95, 96, 97, one would expect $\Delta\Delta G_{rem}$ to promote folding by discouraging large cavity formation. However, our results indicate that $\Delta\Delta G_{rem}$ favors the unfolded state. This suggests that the other entropic effect, polymer-solvent interaction entropy, dominates and counteracts the stabilizing excluded volume contribution. This destabilizing role of interaction entropy is consistent with Olgenblum et al. 54, whose findings showed that the polymer-solvent interaction entropy opposes folding for two miniproteins in various sugar solutions, including $\alpha$-glucose, trehalose, and sucrose.

### v2

@@ 1 Introduction
While this work was motivated by formulation design, the insights of this work are also relevant to understanding the influence of cosolvent environments on macromolecular folding and stability. Given that cytoplasm is a crowded environment comprising proteins, metabolites and osmolytes, it is important to understand protein folding equilibria in crowded environments.64, 65, 66, 67, 68, 69 By quantifying the modulations of polymer conformations from sugar-induced compositional and entropic effects, our framework offers a simplified molecular perspective relevant to crowding phenomena in cellular environments.
@@ 2.1 System Setup and Molecular Dynamics Simulations
We simulated a united-atom HP model with 26 monomeric units, where each monomer represents a methylene ($\mathrm{CH_{2}}$) unit in terms of size and mass. The Lennard-Jones parameters used for the polymer model are the same as those used previously to represent $\mathrm{CH_{2}}$.70, 71, 58 ($\sigma=0.373$ $\mathrm{nm}$ and $\epsilon=0.5856$ $\mathrm{kJ/mol}$). The CP model is similar to the HP model with additional partial charges ($\pm 0.5$ e) assigned to 4 beads (Fig. 1a). The charges were arranged in a (++––) sequence to mimic localized charged patches on protein surfaces, which promote intrachain electrostatic attraction and a hairpin-like folded configuration. The net charge of the CP model is zero, ensuring that any observed differences arise from local electrostatic effects rather than overall charge imbalance. Replica exchange umbrella sampling (REUS) simulations (see Section 2.2) were performed using GROMACS 2021.7 with the PLUMED 2.9.0 patch applied. Kirkwood-Buff derived parameters (KBP) developed by Cloutier et al. 72 were used for the sugars ($\alpha$-glucose, $\beta$-fructose, trehalose, sucrose) considered in this study (Fig. 1b). The KBP models employ an all-atom representation of carbohydrates and more accurately describe protein-excipient-water interactions than standard force fields.
We employed the TIP4P/2005 water model73 to describe water. TIP4P/2005 has been shown to accurately predict several water properties such as density, viscosity, diffusion coefficients, and the hydrogen-bonding network 74, 73. Since the original KBP were developed for use with the TIP3P water model,75 we assessed their applicability to the TIP4P/2005 water model. We calculated the Kirkwood-Buff integral for the sugar-water systems and compared the results with those reported by Cloutier et al.72. The values obtained are within the range of experimental error, confirming the compatibility of KBP with TIP4P/2005 model (see Table S1 for details). We performed MD simulations of HP and CP models in sugar solutions of concentrations ranging from 0 to 2.0 M (see Table S2 for further details). Monosaccharide concentrations were studied up to 2.0 M and disaccharide concentrations were limited to 1.0 M to ensure an equivalent monomeric unit concentration.
@@
The systems in all simulations were first energy minimized using the steepest descent algorithm.76 NVT equilibration was carried out for 1 ns at 300 K, followed by a 1 ns NPT equilibration at 300 K and 1 atm. During equilibration, temperature was controlled with the V-rescale thermostat,77 and pressure was controlled with the Berendsen barostat.78 Following equilibration, NPT production runs were completed using the Nosé-Hoover thermostat79 and Parrinello-Rahman barostat.80 Production runs were completed for 100 ns per window for polymer/sugar/water REUS simulations. 100 ns was sufficient to achieve convergence, as demonstrated by the evolution of the PMF with simulation time (SI Fig. S1). The Particle Mesh Ewald (PME) algorithm was used for long-range electrostatic interactions with a cut-off of 1 $\mathrm{nm}$.81 A reciprocal grid of 42 x 42 x 42 cells was used with $\mathrm{4^{th}}$ order B-spline interpolation. A single cut-off of 1 $\mathrm{nm}$ was used for van der Waals interactions. The neighbor search was performed every 10 steps. Lorentz-Berthelot mixing rules82, 83 were used to calculate non-bonded interactions between different atom types, except for polymer-water oxygen interactions. The polymer-water oxygen interactions were adjusted such that the folded and unfolded states of the polymer in pure water were approximately evenly distributed in straightforward MD simulations.70, 71
@@ 2.2 Replica Exchange Umbrella Sampling
REUS was employed to sample the conformational free energy landscape of the polymer in different sugar solutions and obtain the potential of mean force (PMF) as a function of the polymer radius of gyration ($R_{g}$). 12 evenly spaced umbrella potential window centers from 0.35 to 0.90 nm were used, with a spacing of 0.05 nm between consecutive centers. A harmonic bias potential of 5000 kJ/mol/nm2 was applied to all the windows except the window at 0.45 nm, where 1000 kJ/mol/nm2 was applied. This choice of force constants was adopted from our previous work on a similar system, where it ensured adequate sampling and convergence of the free energy profile.70, 71
@@
where $R_{g,cut}$ is determined as the transition point between the folded and unfolded states where $\frac{dW(R_{g})}{dR_{g}}=0$. $W(R_{g})$ is the PMF, calculated as $W(R_{g})=$} $k_{B}T\ln(P(R_{g}))$}. The Weighted Histogram Analysis Method (WHAM) was used to reweight biased probability distributions obtained from REUS simulations.84
@@ 2.3 Thermodynamic Decomposition
We decomposed the PMF into thermodynamic components to further investigate the role of sugars in HP and CP folding. Following the methods established in prior studies,59, 85, 86 the PMF can be separated into two contributions – intrapolymer interactions and solvent-mediated interactions ($W_{solv}$). The former can be estimated from simulations of polymer in vacuum ($W_{vac}$), leading to the expression:
@@
Here, both $W(R_{g})$ and $W_{vac}(R_{g})$ are obtained directly from simulations: $W(R_{g})$ is the potential of mean force (PMF) for the polymer in solution, and $W_{vac}(R_{g})$ is computed for the same polymer in vacuum. The solvation contribution, $W_{solv}(R_{g})$, is then evaluated from Eq. 2 as their difference.
@@
Consistent with the solvation theory approaches, 87, 88, 89, 90, 91, 92, 93, 94, 42, 95 $\Delta G_{solv}$ can be conceptually decomposed into two steps: (1) the reversible work required to create a polymer-sized cavity in the solvent, and (2) the energy contributions arising from "turning on" polymer-solvent interactions upon solvation. This relationship was expressed by Arsiccio et al.42 as:
@@
Here, $\Delta G_{ex}$ represents the free energy associated with creating a cavity in the solvent (also known as the excluded volume contribution), and $\Delta G_{si}$ represents the contribution from polymer-solvent interactions (also known as the soft-interaction contribution). The cavity formation step primarily reflects the energetic cost of disrupting solvent–solvent interactions and increasing the solvent surface area around the polymer. The small PV-work associated with the change in system volume is included implicitly, but it is negligible at ambient pressure, making the interfacial (surface-tension) contribution dominant.
The total enthalpic contribution ($\Delta H_{ex}$ + $\Delta H_{si}$) can therefore be approximated as the polymer–solvent interaction energy, assuming negligible PV-work:
@@
In summary, $W(R_{g})$ and $W_{vac}(R_{g})$ were obtained directly from simulations of the polymer in solution and vacuum, $W_{solv}(R_{g})$ from their difference, $\Delta E_{pw}$ and $\Delta E_{ps}$ from energy decomposition via gmx rerun, and $\Delta G_{rem}$ by difference using Eq. 8. This approach cleanly isolates solvent-mediated contributions governing sugar-induced polymer stabilization.
@@ 2.4 Preferential Interaction Coefficient
where $\mu$ is the chemical potential, $m$ is the concentration in molality ($\mathrm{mol/kg}$), subscripts $p$, and $s$ refer to polymer and sugar, respectively, and T and P refer to the temperature and pressure. This parameter is calculated in simulations using the two-domain formula36, 96, 97 given by:
@@ 3.1 Effect of Sugars on Polymer Folding
The free energy of unfolding provides a measure of the stability of the ensemble of the folded vs. unfolded states of the polymer. For monosaccharides, the free energy of unfolding increases as the sugar concentration increases up to 1 M. This is consistent with previous observations using the linear extrapolation method, which assumes that the free energy of unfolding varies monotonically with cosolvent concentrations and is commonly used to estimate stability changes in the presence of denaturants or stabilizers. 98, 99, 42 However, at 2 M concentration, the free energy of unfolding decreases, suggesting a non-monotonic, concentration-dependent influence of sugars on polymer stability. This shift occurred at lower concentrations for disaccharides (0.5 M) than monosaccharides, consistent with the trends observed in the unfolded state region of the PMF.
@@ 3.2 Role of Preferential Exclusion on Polymer Stabilization
We compute the preferential interaction coefficient ($\Gamma$) for the HP and CP models in different sugar solutions. Our results consistently show positive $\Gamma$ values across all sugar solutions, indicating preferential accumulation of sugars around HP and CP. The magnitude of $\Gamma$ increases with sugar concentration, with disaccharides exhibiting higher values than monosaccharides at the same molar concentrations. To further assess whether these trends arise purely from molecular concentration or from the total sugar content, we also compared $\Gamma$ values on a mass basis, that is, for solutions containing approximately equivalent total sugar mass (e.g., 1 M trehalose vs. 0.5 M glucose). Even under these mass-equivalent conditions, disaccharides continue to exhibit higher $\Gamma$ values than monosaccharides, with the notable exception of 2 M $\beta$-fructose, which shows values comparable to 1 M sucrose. This comparison is relevant for formulation design, where excipient levels are often specified by weight percent rather than molar concentration
The preferential exclusion mechanism 43, 42 suggests that stabilizing sugars are preferentially excluded from (folded) proteins, resulting in negative $\Gamma$ values. However, our results show positive $\Gamma$ values even in cases where the folded states of HP and CP are thermodynamically favored, seemingly contradicting this mechanism. This apparent inconsistency can be reconciled by recognizing that preferential exclusion is fundamentally governed not by the absolute magnitude of $\Gamma$, but by the difference in $\Gamma$ between the folded and unfolded states. Specifically, the mechanism assumes greater exclusion from the unfolded state, and it is this difference in $\Gamma$ that contributes to the stabilization. To capture this effect, we examine the difference in preferential interactions: $\Delta\Gamma=\Gamma_{folded}-\Gamma_{unfolded}$. At lower sugar concentrations, $\Delta\Gamma$ is positive, indicating stronger sugar accumulation around the folded state. This correlates with an increase in unfolding free energy with concentration, suggesting increasing stabilization of the folded state. At higher concentrations, $\Delta\Gamma$ becomes negative, and this transition coincides with a decrease in unfolding free energy with increasing sugar concentration. This trend is consistent with the Wyman linkage relation 100, which links changes in solute binding to changes in macromolecular stability. Wyman linkage relation for folding/unfolding equilibria can be expressed as:100
@@ 3.3 Thermodynamic Decomposition of Unfolding Free Energy
Figure 5: Decomposition of HP unfolding free energy contributions in (a) $\alpha$-glucose, (b) $\beta$-fructose, (c) trehalose, and (d) sucrose solutions. Positive values make the unfolding in the presence of sugar molecules less favorable than in pure water. Increasing additive concentration is denoted by increased shading (light to dark).
Figure 6: Decomposition of CP unfolding free energy contributions in (a) $\alpha$-glucose, (b) $\beta$-fructose, (c) trehalose, and (d) sucrose solutions. Positive values make the unfolding in the presence of sugar molecules less favorable than in pure water. Increasing additive concentration is denoted by increased shading (light to dark).
@@
For HP in monosaccharide solutions (Fig. 5a, b), polymer-sugar interactions ($\Delta\Delta E_{ps}$) stabilize the folded state up to a concentration of 0.5 M. At 1 M, this effect reduces, and by 2 M, these interactions favor the unfolded state. This transition suggests that at higher sugar concentrations, additional binding sites on the polymer that become accessible upon unfolding interact preferentially with sugars, thereby stabilizing the unfolded conformation. This is also largely supported by the increased $\Gamma$ values for the unfolded state at high sugar concentrations. These findings explain the decrease in $\Delta\Delta G_{u}$ observed at high sugar concentrations. Conversely, polymer-water interactions ($\Delta\Delta E_{pw}$) favor unfolding at concentrations up to 0.5 M but promote folding at higher concentrations. $\Delta\Delta G_{rem}$, representing the entropic effects, consistently favors unfolding in all sugar solutions relative to pure water.
In disaccharide solutions (Fig. 5c, d), a similar trend is observed, but the stabilizing effect of $\Delta\Delta E_{ps}$ shifts toward favoring the unfolded state as early as 0.5 M. This earlier onset of stabilization relative to monosaccharides can be attributed to the larger molecular size of disaccharides: a 0.5 M disaccharide solution contains approximately the same total sugar mass as a 1 M monosaccharide solution. Consequently, trends that appear at 1 M for monosaccharides tend to emerge at roughly 0.5 M in disaccharide solutions. This mass-equivalent comparison helps explain why $\Delta G_{u}$ decreases at lower molar concentrations in disaccharides relative to monosaccharides. While this mass-based comparison captures broad mass-dependent trends, deviations among sugars at equal mass concentrations indicate that other molecular features also contribute to the observed behavior, as explained further in Section 3.4 and 3.5.
@@
As discussed in Section 3.3, $\Delta\Delta G_{rem}$ consists of entropic effects from cavity creation (excluded volume contribution) and polymer-solvent interactions (referred to as the interaction entropy contribution). Previous studies suggest that the excluded volume contribution is directly related to solvent surface tension. 43, 49, 42 Specifically, higher surface tension increases the energy required for cavity formation, favoring smaller cavities and thus stabilizing the folded state. Since all sugars considered here increase the air-water surface tension 43, 101, 102, 103, one would expect $\Delta\Delta G_{rem}$ to promote folding by discouraging large cavity formation. However, our results indicate that $\Delta\Delta G_{rem}$ favors the unfolded state. This suggests that the other entropic effect, polymer-solvent interaction entropy, dominates and counteracts the stabilizing excluded volume contribution. This destabilizing role of interaction entropy is consistent with Olgenblum et al. 54, whose findings showed that the polymer-solvent interaction entropy opposes folding for two miniproteins in various sugar solutions, including $\alpha$-glucose, trehalose, and sucrose.